OpenStudy (anonymous):

please explain how to solve this step by step? 9(8 - x) = 3x.

OpenStudy (anonymous):

Ok, well, first let's use the distributive property on the left hand side. What do you get?

OpenStudy (anonymous):

72x9x?

OpenStudy (anonymous):

Close. 8 and 9 don't have an x in their term, so that would just be 72, making it 72-9x

OpenStudy (anonymous):

so the answer would be what?

OpenStudy (anonymous):

well, you want to solve for x, right? so get the x-terms onto one side

OpenStudy (anonymous):

There are two ways you can solve this. 1. Since 9 and 3 are divisible by 3, divide both sides by 3. This gives you 3(8 - x) =x 2. Then, distribute out the 3 on the left hand side (PEMDAS). This gives you 24 - 3x = x. 3. Get all the x-terms on one side. this gives you 24 = 4x. 4. Divide both sides by 4. This gives you 6 = x The other way is to 1. Distribute out the 9, giving you 72 - 9x = 3x 2. Add 9x to both sides, giving you 72 = 12x 3. Divide both sides by 12. This gives you 6 = x.